OpenStudy (anonymous):

y =3/2 x + 4 –3x + 2y = 8

OpenStudy (anonymous):

@Hero can u help me please

hero (hero):

replace y in the second equation with it's equivalent expression. We know y = (3/2)x + 4 so... -3x + 2((3/2)x + 4) = 8 Which becomes -3x + 3x + 8 = 8

hero (hero):

Oh, that's interesting.

hero (hero):

@reallybubbles

OpenStudy (anonymous):

yes

hero (hero):

Are you able to solve for x?

OpenStudy (anonymous):

solve for x??

hero (hero):

If possible, yes

OpenStudy (anonymous):

how

hero (hero):

Well if you have done the substitution correctly, you should have gotten -3x + 3x + 8 = 8 but... -3x + 3x = 0 and 8 = 8 which means it is not possible to solve for x.

hero (hero):

However, the equation is still true so that must mean infinite solutions for the system.

OpenStudy (anonymous):

thank u